FEL Monthly Highlights
July 2001

Jefferson Lab and the scientific community were saddened this month by the passing of our dear friend and colleague, Nathan Isgur, who lost a courageous battle with cancer.  For the last decade Nathan served as Jefferson Lab’s Chief Scientist, Head of the Nuclear Theory Group and the University Relations Office.  A memorial service is being planned for September in Williamsburg.  Information will be posted on the following website: http://www.jlab.org/Isgur/.

We achieved a new average power record for the FEL during July test runs: 2.1 kW- an increase of 24% over the previous record of 1.7 kW.  The FEL was operated with a wiggler efficiency of 1%. For the record the achieved power level and wiggler efficiency are now twice the original design specification.   In addition we tested FEL operation with the gun operated at the peak charges (135 pC) needed for the 10 kW Upgrade.  Details of both experiments are given in the Operations Section.

Highlights from the start of the summer user run in July include the following:

We analyzed the terahertz radiation emanating from the recirculation bend magnets.  The quantity of terahertz radiation (~watts) will be very useful for experiments and its signature will be a useful measure of the electron bunch characteristics.

The FEL was used to continue Gunter Leupke’s superb series of measurements on the nature of hydrogen defects in silicon.  A second article on this work has been accepted by Physical Review Letters and another article is in preparation for Science.

With respect to the FEL Upgrade project: we sent out the first procurement package for one of the dipole magnets needed for the upgrade; prepared the procurement package for the new drive laser system; nearly completed the Zone 3 RF system; ran the first optical tests in our new mirror test stand; and readied the optical klystron for its magnetic measurements next month.

This month (July) we received funding authorization ($404k) from ONR and DOE for the JTO contract for laser materials damage (i.e., lethality) studies.  We thank all involved at JTO, ONR and the DOE JLab Site Office for blazing the path through the bureaucracies.

We received a summary report from SURA’s Maritime Technology Advisory Committee (MTAC) who reviewed our Navy funded FEL program on June 26.  MTAC graded our plan for a 100 kW Upgrade and felt the plan was credible and made several recommendations for strengthening our relationship with the Navy.

With the cooperation of the Jefferson Lab DOE Site Office we prepared a draft interagency agreement for the AF funded UV FEL project.  We have been notified by AFRL and OSD that the funds for the first phase (FY01-02) effort for this project should be available by the end of the month. (As of press time we are still waiting.)

We have fixed the date for the next upgrade project semiannual review as Nov. 8-9.

Project Cost Performance:
The project (Phase 1) budget for the period June 1, 2000 to September 30, 2001 is $9,029k.  The Phase 2 project was approved for $4,500k for a performance period of February 1, 2001 to September 30, 2002.  The project through the month of July has a total of $4,972 of performance scheduled (assuming the project started at the originally planned start date of April 1, 2000).  The work performed through the end of July was $5,523, which is 61% complete vs. 55% scheduled.  The actual cost accrued through July totals $6,899.  This results in a schedule variance of +$551 and a cost variance of -$1,376.

We are on track for having the first phase of the project fully obligated by September and full costing of all outstanding commitments by Dec. 30.  The total project cost to date-actual costs plus commitments is $8,655k.  By definition the total project cost will not exceed the project budget of $9,029k. We are in the process of closing out the Phase 1 project cost-performance by Sept. 30.   This process includes examination of all outstanding project commitments within the Phase 1 budget and an analysis any possible cost or schedule variances within each WBS element by the project end date.

WBS 3 (Beam Physics):
Two days were spent doing orbit studies to benchmark our model of the lattice and optimize the beam transport for July’s high power test.  Driver issues for MW class performance was documented (JLAB-TN-01-0340).

Three pass beam was circulated through the Demo driver, momentum spread was reduced, and energy recovered beam was observed at the dump.  Pictures (of 3 passes at the end of the module, 2 in the backleg, and 1 at the dump) are available at:
http://www.jlab.org/~douglas/FEL/configuration_updates/3passvg.ppt   A tech note will be forthcoming.

Upgrade octupole specifications were published as JLAB-TN-01-036, which is available online at:

Transverse linac optics for 1 pass CEBAF-ER (energy recovery experiment) were developed and will eventually be documented.

WBS 4 (Injector):
We ran an additional curve on the N2 implanted faraday shielded Ti electrode at a 5 mm gap this time. The performance was worse than 1um polished Ti.  The spacer nipple (10-6 MV/m) for the gun has been manufactured.  We performed a dimensional stack up check on the gun drawings to confirm the proper anode to cathode gap was kept.

The gun chamber drawings have been signed off and are in the shop for fabrication.  The top hat flange assembly drawings were also signed and submitted.

The cathode support tube and the cathode stalk were submitted to the machine shop for fabrication.  The ceramic stack was welded together and is waiting for leak checking, the top hat flange being manufactured is needed for this.

The GaAs wafer in the existing gun passed the 4k Coulomb delivery mark this month.

Gun HVPS - Progress is being made with the design of the new Gun Tank.  It will be a cylinder 70" in diameter and will not require any removal of the short concrete wall or pad.  All of the parts of the HVPS returned from Glassman are being identified for their mounting in the tank.

WBS 5 (SRF):
Receipt inspection of the first vacuum vessel and space frame continues.  Minor discrepancies have been noted including some marks on o-ring sealing surfaces and general cleanliness.

Project engineers visited the End Can and Thermal Shield vendors.  Progress continues on cryomodule component schedule with no significant technical issues.

WBS 6 (RF):
Zone 3 - The HVPS for this zone was installed and is being connected.  The circulators, waveguides, couplers, shorts and test couplers were also installed this month.  Connection to AC power will wait until after the summer run is over.  Installed 4 RF Control Modules with 4 test units due this month.  All 8 klystrons, Mod Anode, and Filament boards were installed.  LCW and Instrument Air were activated for this zone.  Power and control cables are being wired between the HVPS and the High Power Amplifier (HPA).  Control Power is connected to the HVPS.  The 5 watt amplifiers and their power supply have not yet been installed.

Zone 4 - Waiting for 7 RF Control modules, the installation of the last circulator, and the completion of the punch list.

Quarter Klystrons - September 12th is the latest date to witness Factory Tests of the 100 kW klystrons at CPI.  The carts for the klystrons are being built in the machine shop.

Quarter HVPS - Contracts were placed for the new Transformers and Inductors with NRL.  The Rectifier Assembly was ordered from CKE.

WBS 8 (Instrumentation):
The 3"  beam current cavity design is complete and an order is being placed for 5 pieces. The pulse truncator was modified and is awaiting further testing. This device allows for 0.5 microsecond increments in the pulse width for the FEL laser pulse. A prototype of a new BPM electronics is being worked on, testing should start in a couple weeks. Thanks to the EID group for there continued support.

Prototype boards were ordered for the new VME Timing Board. Cables were pulled and the new LINAC Patch Panels installed.  James built two new Pyro-electric Beam Position Monitors for testing. Two user cameras were repaired, and 6 more ordered.  Dan built two Temperature Monitoring Chassis for RTD readout.

Patrick has the filter wheel working using the Rabbit TCP/IP interface and says the Optical Power Meter is coming along fine.  This will allow for rapid connection of serial devices to the network and remotely monitor and control those devices without the FEL Users logging into EPICs.

Mike updated the search feature of the task system by offering a choice of the last 10 entries made for a particular task, or listing all entries pertaining to that task.  He also added the VME card inventory to Laser and condensed some databases into a more useable format.

A new intercom system is being installed in each of the user labs, the OCR and the Drive Laser Clean Room.  This should aid in communications and reduce noise in the control room from ringing phones.

Absolute value/Sample and Hold boards for the system have been designed and prototype boards are on their way.  An associated 32 channel Instrument Amplifier board has also been designed and the artwork completed.  A programmable delay card for the new pulse controller has been designing has been built and is in the final stages of testing.  Several VME GPVS cards were successfully tested as well.

A 32 channel sample and hold buffer board has been designed.  The design is being reviewed and prototypes will be out soon.  This board will work with the OBPM system, however it will also have applications in future systems as well.

The programmable delay card has been successfully tested.  There are still some programming issues in EPICS to be resolved.  This should prove to be a very versatile system once all items have been completed.

The task system on http://laser has been updated grouping tasks by system for ease of locating specific tasks and historical archiving.  The website has also been updated to include a cable database.  Both of these items will continue to be upgraded as more information becomes available.  A new feature documenting all tasks associated with each system has been added to the website.  Currently only I & C and Optics tasks are available; however, the system is available for all groups associated with the FEL to contribute to.

Drawing updates continue with several on hold pending final design changes.
Drawings completed:  OBPM/Picomotor Test Box Wiring, Assembly and Fabrication; Programmable Delay Card Schematic, Fabrication and Artwork; Pyro-electric Detector Buffer/ Amplifier Schematic, Fabrication, Assembly and Artwork and the revision of the Beamviewer Solenoid Box Wiring Diagram.

Check Plots Received:  Picomotor Relay Wiring Diagram, Molectron Quad OBPM Amplifier Schematic, Fabrication Assembly and Artwork and Absolute Value/Sample & Hold Board Schematic, Assembly, Fabrication and Artwork.

Drawings at EECAD:  Analog/Differential Driver Board Schematic; Pyro-Electric Detector Buffer Amplifier Board Schematic, Assembly, Fabrication and Artwork; 32 Channel Sample and Hold Board Schematic, Assembly, Fabrication and Artwork; Ion Pump Power Supply Controller Schematic; User Lab Shutter Control Box Wiring, Schematic, Assembly and Fabrication; MPS System Drawing; Generic Temperature Monitoring Chassis Schematic, Wiring, Fabrication and Assembly.

We regretfully say goodbye to James Jones as he leaves us for the Baltimore area.  He will remain on as a casual employee for the time being.  Good luck James.

WBS 9 (Transport):
Optical Chicane Dipoles (GW)
o We will initiate a separate procurement for these magnets now and not wait for the completion of the design of
    the 180-degree dipoles (GY) as originally planned.  This package is complete while the GY requires some
    final adjustments to the model followed by the final tweaks to many of the details.
o The drawings and the coil, core (and parts) and the assembly specifications for the GW were written and
    signed.  Those specifications will serve as a template for the future dipole procurements.
o We wrote the requisition and are starting discussions with Procurement to insure a smooth procurement
    process for the several batches of dipoles we will need.
o At the end of this period, bid documents are ready to be sent out by Procurement.
Injector Dipoles (DU/DV)
o The DULY continues working on the revisions to the drawings of the Injector Dipoles (GU & GV). These
    revisions are going through internal review at DULY and have started to be sent to us for a final check.
Arc Dipoles (GY, GX, GQ)
o AES continued design of the Reverse Bend (GQ). Coil leads, hoses, manifolds and cover layout and details as
    well as the field clamps were completed and the position adjusters placed and assembly refinement started.
o In magnetic modeling, the final configuration of the Bend Dipole (GX) was achieved that fulfilled David Douglas’
    specifications.  The bulk field was achieved over 67% of the magnetic length to one part in 10,000 (This
    worked in the IR Demo) and the current to do this will serve as the anchor current for the Dipole strings.
    The effective length was achieved, shortening the iron length by 1 cm.
o Additional manipulation of the magnetic model of the GX showed that moving the field clamps by 3/8 inch is
    the equivalent of a 0.5 change in iron length.
o In magnetic modeling, they are now revisiting the 180 Dipole (GY).  It is going through final adjustment that
    hinged on the design of the GX.  The bulk field and current is being matched to the GX by adjusting the Purcell
    gap and the gap. Following that, AES is checking the saturation due to our reduced field path length
    correctors and will then adjust the end face position and the angle of the path length coil slot to bring the field
    integrals throughout the good field region into 180 degrees of effective length.
QX (3 inch quad)
o We made some progress and gained insight in eliminating higher order multipoles due the pole tips’ end field
    roll off.  We brought the sum of the multipoles on the Prototype from 0.62 % to 0.42 with more progress
    expected. This will be a last tweak to the core drawings that we can introduce after the core order is placed.
o The core contract was placed with New England Techni-Coil, a quality vendor who has supplied us in the past
    and we are sending them steel to start core production.
o Magnet Enterprises International is perfecting their winding technique on the first article coil.
QT (Trim Quad)
o The magnetic model is acceptable.  Most of the design and details are drawn.
o The interface of the magnet to the measurement stand and the details of the advantageous use of the known
    dimension conductor that are being worked.
o All the pieces of the 3D drawing model were modified to the narrow configuration as a precursor to starting on
    the 2D detail drawings.
o Robin Wines has not been able to continue with magnetic modeling over the last reporting period because of
    the octupole work below, her checking of the magnetic quality, specifications and drawings of the GW dipole
o The task order for DULY Research to start magnetic modeling and design was sent to Procurement.
o DULY Research responded to the task order for magnetic modeling and design.  In response, we are refining
    our specs and reducing the task to magnetic analysis until we understand more about the final configuration.
    This will allow DULY to get a better picture and be more realistic with regard to the drafting design task.
o Dave Douglas refined his specification for the octupole in a published tech note.  The final result is a further
    reduction in field by over half but higher quality, going from 10% to 1%.
Beam Line and Vacuum
o The review period for the set of the drawing of the girders for the return leg was completed.
o Work continued on the drawing set for the girders between Cryomodules.
o Work started on the layout of the injection and recombination area.
o Work continued on the Interface Control Document for the task definition for the Arc and Optical Chicane
    chambers.  The work concentrated on the specialty chambers through which the optical beam has to traverse.
    Better definition of the optical requirements lead to less drastic requirements on the specialty chambers through
    which the optical beam traverses.
o The Dipole Steel order was placed with Bethlehem Lukens Steel with a 10-week delivery.
o We generated the first pass at the paperwork for the CRADA for the collaboration between Jefferson Lab and
    STI Optronics to test out their compact 180-degree bend design on the FEL Upgrade machine.

WBS 10 (Wiggler):
Analysis of dispersion correction coils around the entire magnet and around just the center pole indicates that the field error arises from a 3D solenoid component caused by excess reluctance on the end poles.  This is best compensated with a solenoidal coil around the whole magnet.  The dispersion section strength is unchanged when this coil is used while the coil around a single pole predominantly changes the dispersion section strength and changes the steering as a second order effect. We have therefore settled on a design for a 3 turn correction coil around the entire magnet.  Michael Necaise is working on this.

These corrector coils were wired up and the wiggler is now ready for measurements.  The measurement apparatus is completed and is in the process of being calibrated before measurements can start.  That is expected to happen in August.

We hope to have the dispersion section chamber ready for the final measurements of the dispersion section.  Final welding was started on the dispersion section vacuum chamber.  There is one machine step after that and the chamber will be ready for installation in the dispersion section magnet for measurements.

WBS 11 (Optics):
Detailing continues on the internal mounting hardware in the optical cavity chambers.  We received drawings of the mirror design employed by the AVLIS, and now NIF (courtesy of R. Chow - LLNL), and are evaluating their applicability.

We decided to abandon (at least for now) use of two insertable viewers to route beam to the diagnostics mounted between the outcoupler chamber and the collimator.  The designer found that dual-action air cylinders exist and should do the job.

Our laser interferometer was installed by the vendor's application engineer.  This will be used in general to confirm optics' specifications, and specifically, with the mirror test stand.

The new drive laser specifications and objectives were turned over to Procurement.

We are detailing the internal hardware in the optical cavity chambers, specifically: (1) edge-cooled mirror holders for the outcouplers, and the modification to the mirror mounts to accommodate them (2) y-motion of the entire mirror frame.  The outcoupler mirror holder has been refined sufficiently that it fits within our space constraints.  We've also determined that the vendor for the rotary feedthrough can provide us with a gear ratio that gives us good resolution, and reasonably rapid translation.  The position-sensing portion of the mirror metrology system is being tested; an engineer has been assigned to the team to assist in the design of some of its hardware.

We completed preparations for our first use of the mirror test stand (MTS) to measure the absorption of the 6 micron high reflectors.  Thanks to I & C for installing the RTDs and wiring the feedthrough to allow monitoring while under vacuum. We did our first measurements of absorption levels in our 6 micron HR mirrors in the mirror test stand (MTS).  Data analysis remains to be done, but IR thermometry results are consistent with a loss less than 500 ppm.  Inspected and passed half of our new 1.6 microns optics.

We have made a decision to purchase the UHV-compatible actuators for the mirrors, and a PR was placed by July 27th.

The purchase requisition for the new drive laser has been placed and the RFP is due to be sent out today.

Ops ran smoothly over the first week of the July-August summer run. The first part of the week was devoted measuring ablation at 2.9 microns.   That was followed by X-ray Thompson scattering measurements.  The X-ray measurements successfully mapped the distribution of scattering of 3 micron light for the first time in preparation for a double scattering experiment to be performed in the fall.  That effort was followed by a high power demonstration run.  We were repeatedly able to run to 2.1 kW using 65 pC for 0.97% extraction efficiency, a 24% improvement over our previous CW efficiency.

We then re-cesiated the cathode, this being the first re-cesiation since Feb.16.   The QE improved from 0.5% to 1.9% and we again verified the high power FEL performance after the recesiation.  Since average current was still limited to 5 mA no significant additional power was produced.  We did not take the additional time it would take to optimize the setup for high charge running at 135 pC although we were able to lase strongly there with 1% pulsed efficiency.  That effort was followed by a switch over to 6 micron mirrors and we brought up the system and established lasing for user tests at 5.8 microns for next week.

We also were able to see and measure the spectrum of the collective far IR radiation emitted from the electron bunches in going around our bends.  It is expected that this radiation will be useful as a beam diagnostic and a research tool in its own right.

Three runs in an entirely parasitic mode were accomplished for the nanotube collaboration (CWM, NASA, PSU).  They are continuing to collect nanotubes at various target/laser configurations to optimize the
production rate of single wall tubes.

We operated the FEL the second run week for various users.  Work at the beginning of the week involved testing of polymer surface modification on a resonant absorption line.  We were able to run short macropulses of 4 microseconds at 20 kHz to maximize energy transfer rate.   Parasitically we were able to run X-ray measurements looking for the low energy X-rays produced while lasing around 6 microns.

We then spent two days successfully making measurements on non-linear absorption in dielectrics.  All the samples available were exposed and brought off for analysis.  We then performed alignment tests for near field IR microscopy and performed saturation tests for Beam Position Monitor electronics.   We installed a calibrated pyroelectric detector and were able to measure the total collective edge radiation power (~0.2 W) in the THz regime through a 1 cm diamond window. Since this is produced in short pulses and is a very intense tunable source at this wavelength it opens the possibility of a number of future scientific tests.  X-ray measurements continued through the week but struggled with signal level due to the high background and large window absorption at the 3 keV energy range. The final day was spent measuring absorption in 6 micron mirror
coatings.  Preliminary results show the losses are very small, less than 100 PPM.

For the third run week we provided support to G. Luepke (College of William and Mary).  A beautiful set of data was obtained on H-lifetime vs. temperature from 77k to RT.  This continues Prof. Luepke’s string of publication quality measurements every time he runs.

At the beginning of the fourth run week we re-established and stabilized third harmonic lasing at 1 micron for the rest of the week's tests.  We then moved on to utilization of the 1 micron light.  Tuesday was spent testing the ability of the 1 micron pulses to machine copper coated steel in a trepanning setup.  We also tried to ablate solar sail material and some cutting of glass.  On Wednesday we attempted to produce carbon nanotubes with 1 micron light in the same manner where we had been successful at 6 microns.  However in this case no production was seen.  We spent the rest of the week producing harmonics of the 1 micron light in doubling and quadrupling crystals and using that light for exposure of photosensitive glasses with our collaborators from Aerospace Corp.  The FEL operated stably and reproducibly throughout this period.